y=sec^(−1) (x) ⇔ secy=x,f or y∈[0,π/2)∪[π,3π/2).


With this choice of principal branch the derivative of sec^(−1) (x) is


d/dx*sec^-1 (x)= 1/x*√(x^2 -1)


However, as mentioned in the side remark in Section 4.9, there is another common

definition of sec^−1(x), which is y=sec^(−1) (x), which is


y=sec^−1⁡(x) ⇔ sec⁡y=x, for y∈[0,π/2)∪(π/2,π].


Show that, with this choice of principal branch to define the inverse secant function, its derivative becomes


d/dx*sec^(−1) (x)=1/|x|*√(x^2 -1)




1
Expert's answer
2021-10-22T01:01:21-0400

"secy=x"

"(secy)'=y'secytany=1"


"y'=\frac{1}{secytany}"


"tany=\sqrt{sec^2y-1}=\sqrt{x^2-1}"


"\frac{d}{dx}sec^{-1}x=\frac{1}{x\sqrt{x^2-1}}"